Mrs. Kwok is 6 years more than 3 times as old as her daughter. Eight years ago she was 2 years less than 9 times as old as her daughter was then. Find each of their ages.
Answer by MissionPossible(18)

Mrs. Kwok's age = K
Her daughter's age = D
Mrs. Kwok is 6 years more than 3 times as old as her daughter.
K - 6 = 3D -------(1)
Eight years ago she was 2 years less than 9 times as old as her daughter was
(K-8)+ 2 = 9(D-8) --------(2)
K - 6 = 9D - 72 ---------(3)
K - 6 = 9D - 72 ---------(3)
K - 6 = 3D ---------(1)

Eq(3) substract Eq(1)
0 = 6D - 72
6D = 72
D = 72/6 = 12
Eq(1) K - 6 = 3D ---------(1)
K - 6 = 3*12
K - 6 = 36
K = 36 + 6
K = 42
Mrs. Kwok's age = K = 42 years
Her daughter's age = D = 12 years Ans.